13. (3pts) An engineer wishes to investigate the effect of adhesive type, conductor material, cure time, and types ot post coatings on bond strength when mounting an integrated circuit (IC) on a certain substrate. Two adhesive types, two conductor materials, three cure times, and four types of IC post coatings are under consideration. H many ways can these four factors be chosen?
Answer #1

total number of choices =N(number of adhesive type)*N(type of conductor material)*N(type of cure time)*N(type of post coatings)=2*2*3*4=48